OpenStudy (anonymous):

Kevin picked 13 more apples than Andy If Kevin picked 30 apples how many did Andy pick?

OpenStudy (anonymous):

Need the equation.. Let a = apples Andy picked.

OpenStudy (anonymous):

I know he picked up andy picked 17

jimthompson5910 (jim_thompson5910):

good, what did you need help with?

OpenStudy (anonymous):

Just need an equation on how to solve

jimthompson5910 (jim_thompson5910):

ah gotcha

jimthompson5910 (jim_thompson5910):

if a = apples Andy picked and "Kevin picked 13 more apples than Andy", then Andy's Apple Count +13 = Kevins apple count a+13 = 30 So the equation is a+13 = 30

OpenStudy (anonymous):

^Thanks Buddy :)

jimthompson5910 (jim_thompson5910):

You'll notice that solving a+13 = 30 will give you a = 17, which is what you previously found. You're welcome.